OpenStudy (anonymous):

NEED HELP! Find the vertex, focus, directrix, and focal width of the parabola. negative 1 divided by 16 times x squared = y Vertex: (0, 0); Focus: (0, -4); Directrix: y = 4; Focal width: 16 Vertex: (0, 0); Focus: (-8, 0); Directrix: x = 4; Focal width: 64 Vertex: (0, 0); Focus: (0, 4); Directrix: y = -4; Focal width: 4 Vertex: (0, 0); Focus: (0, -4); Directrix: y = 4; Focal width: 64

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it would be easier if you wrote it as \(-16y=x^2\)

OpenStudy (misty1212):

then it looks more like \[4py=x^2\]

OpenStudy (anonymous):

Ah, sorry

OpenStudy (misty1212):

yeah i know, but if you multiply both sides by \(-16\) you get \[-16y=x^2\] really we have no work to do here, because the focal width is \(|-16|=16\) and you only have one choice with that answer

OpenStudy (anonymous):

THANKS! @misty1212 would u mind looking at this one please? Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -3 + 3 sin θ y-axis only Origin only x-axis only No symmetry